Assume that if we pick a person at random from a very large group and determine the day of the
week on which they were born, there is an equal chance that it is each possibility (Monday, Tuesday,
etc.). Suppose seven people are chosen from the group at random. What is the chance that no two
persons were born on the same day of the week?
Answer #1

chance that no two
persons were born on the same day of the week =P(all 7 people are born on different days of week)

=7P7 /77 =7!/77 =5040/823543 =720/117649